Please help me solve this equation:
2(x-6)+8=4(x+2)-6
Solution x=

Answer by stanbon(48568)

2(x-6)+8=4(x+2)-6
2x-12+8 = 4x+8-6
2x-4 = 4x+2
2x = -6
x = -3
===========
Cheers,
Stan H.